OpenStudy (anonymous):

(Calc Q > Derivatives > Chain Rule > Double chain rule scenarios?!) So I'm working on some problems for my homework and come across a problem that requires me to apply the chain rule twice. I have no clue where to begin, even though I do understand how to apply the chain rule. The problem in question is: g(x) = 2tan^3(4πx). The answer is 24π(tan^2(4πx))(sec^2(4πx). However having the answer helps me none since I do not know how to apply the chain rule twice... Help?! (MUCH thanks in advance!!!)

OpenStudy (amistre64):

it might be good to rewrite the trig \[ 2~[\tan(4πx)]^3\] it helps to see the chaining parts \[ 2*3~[\tan(4πx)]^2*\tan'(4πx)\] \[ 2*3~[\tan(4πx)]^2*\sec^2(4πx)*4πx'\]

OpenStudy (amistre64):

the chain rule is just a matter of popping out the derivative of an inner function, like a set of russian dolls worry about the outer most function, before focusing on the inner functions

OpenStudy (amistre64):

\[ 2~tan^3(4πx)~:~let~a=tan(4\pi x)\] \[ 2~a^3\to~2*3a^2~a'\] since a = tan(4pi x) , let b = 4pi x a = tan(b) a' = sec^2(b) * b' \[ 2~a^3\to~2*3a^2~a'\] \[2*3a^2~a'\to~2*3a^2~sec^2(b) * b'\] so, what is b'? etc ....

OpenStudy (anonymous):

i still did'n get it... im going to ask simpler question what is the derivative of [2xcos(x ^{2})\]

OpenStudy (amistre64):

show me how you would try to work it so i can see if you are introducing an error in the process

OpenStudy (anonymous):

you mentioned that derivate of 2a^3 is 2*3a*a' how do you get a' from derivative?

OpenStudy (amistre64):

thats what the chain rule does, it pops out a derivative of an inner function right? You did say: " I do understand how to apply the chain rule"

OpenStudy (amistre64):

the derivative of f(a) = f'(a) * a'

OpenStudy (amistre64):

one issue with learning this is that you are not told that the chain rule has always been present from the start of you derivative rules ....

OpenStudy (amistre64):

we know that the derivative wrtx of y is y'; or dy/dx the derivative wrtx of x is simply x'; or dx/dx which equals 1 the derivative of a wrtx is da/dx, or simply a'

OpenStudy (amistre64):

2tan^3(4πx) 2 (a)^3 the derivative is 6 a^2 a'

OpenStudy (amistre64):

the chain rule pops out a derivative of the inside function

OpenStudy (amistre64):

in general\[[f(g(h(n(m(...))))))]'=f'(g)*g'(h)*h'(n)*n'(m)*m'(...)~*~...\]

OpenStudy (anonymous):

plz check my answer... 2x cos (x^2) assume : a= 2x a'=2 b= cos(c) b'=-sin(c) c=x^2 c'=2x derivative: a'=2 b' = 2(-sin(c)) = -2 sin(x^2) c' = -2 sin(x^2)*2x then the answer is -4x sin(x^2)..?

OpenStudy (amistre64):

dont forget that you started out with a product ... 2x cos (x^2) the product rule gives us 2x' cos (x^2) + 2x cos'(x^2) the leftside is simple enough since x' = 1 2cos (x^2) + 2x cos'(x^2) ^^^ inner function pops out a derivative afterwards 2cos (x^2) - 2x sin(x^2) * (x^2)' 2cos (x^2) - 2x sin(x^2) * (2x) 2cos (x^2) - 4x^2 sin(x^2)

OpenStudy (amistre64):

how you organize the results will depend on what best suits you :)

OpenStudy (anonymous):

ok2... i get it now... that's mean my second question is not double chain rule...

OpenStudy (amistre64):

right, it can be expressed as a single popout :)